It takes a groundskeeper 45 minutes to prepare a softball field for a game. If it takes his assistant 75 minutes to prepare the same field, how long will it take if they work together?
Answer by richwmiller(17219)

1/a+1/b=1/x
1/45+1/75=1/x
75/3375+45/3375=1/x
120/3375=1/x
x=3375/120
x=28.125 minutes working together
check
x/a+x/b=1
x/45+x/75=1
28.125/45+28.125/75=1
0.625+0.375=1
1.0=1
ok
codewt